# The contact process on dynamical random trees with degree dependence

### v1

@@ 7.4 Transmitting the infection from one star to another
$\begin{split}\mathbb{P}_{\mathcal{T},N}(B_{m}^{1}\mid\mathcal{F}_{t_{m},S})\mathds{1}_{D_{m}}&\geq\mathbb{P}_{\mathcal{T},N}(B_{m}^{1}\mid\mathcal{F}_{t_{m},S}\,,\,\rho\notin\mathbf{C}^{*-x}_{t_{m}})\mathds{1}_{D_{m}}\\ &\geq\frac{\lambda\lfloor\delta\lambda TNp(N,L)\rfloor T}{(\lambda\lfloor\delta\lambda TNp(N,L)\rfloor+1)T+1}\mathds{1}_{D_{m}}.\end{split}$
@@
where in the last step we treated the first time interval $[0,t_{2}]$ differently, since we know that at time 00 the root $\rho$ is infected, and thus we do not need to reinfect it in the first time period. Recalling that we chose $m=\lfloor S(8r+4T)^{-1}\rfloor$, we can conclude
@@
$\leq\mathbb{P}_{\mathcal{T},N}\bigg{(}\bigcap_{i=1}^{m}A_{i}\cap\bigcap_{j=2}^{m}D_{j}\Big{|}\mathcal{F}_{0,S}\bigg{)}$
$\leq\bigg{(}1-b_{\lambda}(N)C_{p}(N)\bigg{(}\frac{\lambda\nu_{1}\kappa_{1}c_{p}}{\lambda+\lambda\nu_{1}+\nu_{1}+1}\bigg{)}^{r}\bigg{)}^{\lfloor S(8r+4T)^{-1}\rfloor},$
@@
We recall that the first claim of Theorem 3.4 is that if $\alpha\in[0,1)$, $\eta\in(-\infty,\tfrac{1-\alpha}{2})$ and (3.7) holds then $\lambda_{2}=0$. In other words, for every $\lambda>0$ local survival takes place with positive probability. The second statement is that if $\alpha\in[0,1)$ and (3.8) holds then $\lambda_{2}<\infty$, i.e. for $\lambda$ large enough local survival takes place with positive probability. Note that it is sufficient to consider $\eta\geq 0$ in order to show this claim because the case $\eta<0$ is covered by the first claim. Therefore, we need to consider two parameter regimes. First let
@@
where $C$ is the universal constant given in Lemma 7.3. Note that if we show the first statement for $\lambda\in\big{(}0,\tfrac{1}{2}\big{)}$ it follows by monotonicity for all $\lambda>0$. The second parameter regime is the following
$\alpha\in[0,1),\,\eta\geq 0\,\text{ and }\,N\in\mathbb{N}\,\text{ such that }\,\lambda=\lambda_{N}=(3T_{N})^{-1}\text{ and }N>C\vee L.$ (R’)
Observe that in both cases we have that $12\lambda T_{N}<$ 1 for all $N$ large enough, which is necessary for several results in the previous sections.
We recall from Section 3.2 that our strategy is to apply [18, Lemma 6.2] to the function
$g(t):=\mathbb{P}^{\{\rho\}}_{N}(\rho\in\mathbf{C}_{t}),\qquad t\geq 0.$
In order to do so, we need to find a non-decreasing function $H$ on $[0,\infty)$ such that $H(x)\geq x$ in a neighbourhood of 00 and show that $g$ and $H$ satisfy
@@
where $S$ is chosen as in (7.18). For simplicity of exposition, we split the proof of Theorem 3.4 into several lemmas. The first lemma gives us a good candidate for the function $H$.
@@
Then for $N$ large enough there exists a constant $c\in(0,1]$ and $\epsilon>0$ such that the function
@@
$p_{in}:=\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}\big{(}x\in\mathbf{C}_{s}\ \text{for some}\ s\in[0,S]\mid\mathcal{F}_{0,S}(x_{1})\big{)}.$
Recall that the process $(\mathbf{C}_{t}^{*-x})_{t\geq 0}$ is defined via the same graphical representation as $(\mathbf{C}_{t})_{t\geq 0}$ but the infection is restricted to $\mathcal{V}^{*-x}=\mathcal{N}_{\rho}\cup\{\rho,x_{1},x_{2},\dots,x_{r-1},x\}$. We begin by noting that due to the monotonicity of the CPDG, we have
$p_{in}\geq\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}\big{(}x\in\mathbf{C}_{s}^{*-x}\ \text{for some}\ s\in[0,S]\mid\mathcal{F}_{0,S}(x_{1})\big{)}.$
Now, Lemma 7.10 tells us that
$\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}\big{(}x\in\mathbf{C}_{s}^{*-x}\ \text{for some}\ s\in[0,S]\mid\mathcal{F}_{0,S}(x_{1})\big{)}\geq 1-(F_{\lambda}(N,r)+R_{\lambda}(N)),$
on the event $\mathcal{S}_{\rho,N}(x_{1})$, where $S$ is defined as in Lemma 7.9, the functions $R_{\lambda}(N)$ and $F_{\lambda}(N,r)$ are defined in (7.19) and (7.20), respectively. Then, using the inequality $1-x\leq e^{-x}$, observe that ($\star$7.11) forces $F_{\lambda}(N,r)$ to be at most $e^{-4}$, i.e.
@@
Before we proceed we need to introduce some further notation. Given a realisation of the underlying tree $\mathcal{T}$ we denote, analogously to (7.5), by $\mathcal{S}_{x,N}^{t-S}$ the event that $x\in\mathbf{P}_{r}$ is a stable star starting from time $t-S$, where $t\geq S$. Furthermore, we explicitly exclude the neighbour in generation $r-1$, but we do omit this in the notation to avoid clutter. In the following, we will condition on the event
@@
i.e. that $x\in\mathbf{C}_{t-S}$ for some $x\in\mathbf{P}_{r}$ which is a stable star starting at time $t-S$. We deduce for $t>S$
@@
Then, by using the fact that $\mathbf{C}_{s}^{*-x}\subset\mathbf{C}_{s}$ for all $s\geq 0$, followed by the tower property and (7.25), we deduce
$\mathbb{E}_{N}\big{[}M_{r}^{S}\big{]}$ $\geq\mathbb{E}_{N}\left[\sum_{x\in\mathbf{P}_{r}}\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}(x\in\mathbf{C}_{s}^{*-x}\ \text{for some}\ s\in[0,S]\mid\mathcal{F}_{0,S}(x_{1}))\mathds{1}_{\mathcal{S}_{\rho,N}(x_{1})}\right]$
$\geq b_{1}\mathbb{E}_{N}\bigg{[}\sum_{x\in\mathbf{P}_{r}}\mathds{1}_{\mathcal{S}_{\rho,N}(x_{1})}\bigg{]}.$
@@
$\mathbb{E}_{N}\bigg{[}\sum_{x\in\mathbf{P}_{r}}\mathds{1}_{{\mathcal{S}}_{\rho,N}(x_{1})}\bigg{]}$ $\geq\mathbb{E}_{N}\big{[}|\mathbf{P}_{r}|\big{]}\mathbb{P}_{N}(\overline{\mathcal{S}}_{\rho,N})\geq\mathbb{P}(\zeta<L)N\mu^{r-1}_{L}\mathbb{P}(\zeta=N)\mathbb{P}_{N}(\overline{\mathcal{S}}_{\rho,N}),$
where we used that $\overline{\mathcal{S}}_{\rho,N}\subset\mathcal{S}_{\rho,N}(x_{1})$ as was pointed out at the end of Section 7.1. In addition, by Lemma 7.1 we have that $\mathbb{P}_{N}(\overline{\mathcal{S}}_{\rho,N})\geq 1-e^{-c_{L}Np(N,L)}$. Thus there exists $b_{2}>0$ such that for $N$ large enough $\mathbb{P}_{N}(\overline{\mathcal{S}}_{\rho,N})\geq b_{2}$. In other words, we deduce
@@
Now, ignore all the infections of $x\in\mathbf{P}_{r}$ by its parent except the first infection. Then, the contact process on the subtrees rooted at vertices $x\in\mathbf{P}_{r}$ which are infected at some time $s<S$ will evolve independently from time $s$ to time $t-S$, since by the branching property of the BGW tree and using that the background process is stationary. Then the vertex $x$ will be infected at time $t-S$ with probability at least $\chi(t)$.
Furthermore, the event that $x\in\mathbf{P}_{r}$ is a stable star from $t-S$ onwards, i.e. that $\mathcal{S}_{x,N}^{t-S}$ is satisfied, is independent of everything that happened before time $t-S$ and by symmetry it follows that $b_{2}$ is a lower bound for the probability of $x\in\mathbf{P}_{r}$ being a stable star for every realisation $\mathcal{T}$, i.e. $\mathbb{P}_{\mathcal{T},N}(\mathcal{S}^{t-S}_{x,N})\geq\mathbb{P}_{N}(\overline{\mathcal{S}}_{\rho,N})\geq b_{2}$.
If we denote by $\widetilde{M}_{r}^{t-S}$ the random number of vertices in $\mathbf{P}_{r}$ which are infected at time $t-S$ and are stable stars starting at $t-S$, we can conclude that $\widetilde{M}_{r}^{t-S}$ stochastically dominates a random variable $M_{\chi}$ that has distribution $\text{Bin}(M_{r}^{S},b_{2}\chi(t))$. Appealing to Lemma 2.3 in [18], there exists $\epsilon_{1}>0$ such that
@@
For that purpose we need to introduce a bit more notation. Let $\mathcal{F}_{s}$ be the $\sigma$-algebra which is generated by the graphical construction up to time $s$. Furthermore, we denote by $\mathcal{F}^{r}$ the $\sigma$-algebra generated by $\Delta^{\text{rec}}_{z}$ and $\Delta^{\text{up}}_{\{y,z\}}$ for all $z\in\mathcal{V}_{r+1}$, i.e. vertices in generation $r$, and $\{y,z\}\in\mathcal{E}$ with $d(\rho,y)=r$, i.e. edges between generation $r$ and $r+1$. In words this means that $\mathcal{F}^{r}$ contains all the information about recoveries in generation $r+1$ and update events between generation $r$ and $r+1$ for the whole time period 00 to $\infty$. Finally we define the $\sigma$-algebra $\mathcal{F}_{s}^{r}:=\sigma(\mathcal{F}_{s}\cup\mathcal{F}^{r})$ for $s\geq 0$.
@@
Recall that on the event $A_{t-S}$ we know that there exists an $x\in\mathbf{P}_{r}$ which is a stable star from $t-S$ onwards and $x\in\mathbf{C}_{t-S}$, i.e. $x$ is infected at time $t-S$. This means we are again in a situation that we can consider the restriction of the graphical representation to a path of length $r$ between $\rho$ and the neighbourhood of $x$. Then, it follows analogously as in the proof of Lemma 7.10 and (7.25) that
@@
where we used that the path is between generation 00 and $r$, and thus independent of everything in generation $r+1$ and higher. Therefore, we can conclude that $h(t)\geq b_{1}$.
@@
Note that on $\mathcal{B}_{\tau}$ we have at least $\delta\lambda TNp(N,L)$ good neighbours in $[\tau+4T,t-1]$. Further, define $\mathcal{I}$ to be the event that at least one of the $\delta\lambda TNp(N,L)$ neighbours which are infected at time $t-1$ infects $\rho$ at a time in $[t-1,t]$ before this neighbour recovers and
@@
Now we find a lower bound for the two conditional probabilities on the right-hand side of the previous equality. First note that $\mathbb{P}(\mathcal{R}_{\rho}\ |\mathcal{F}_{t-1})\geq e^{-1}$. Recall that $\mathcal{N}_{\rho}$ are the neighbours of $\rho$ with degree less than $L$, i.e. $D_{x_{i}}\leq L$ for all $i\in\{1,\dots,\tilde{N}\}$. Denote $\tilde{N}:=\lfloor\delta\lambda TNp(N,L)\rfloor$ and by $x_{1},\dots,x_{\tilde{N}}\in\mathcal{N}_{\rho}$ the $\tilde{N}$ neighbours that are infected at time $t-1$. Note that we can rewrite the event $\mathcal{I}$ as follows
@@
where we used that $\mathcal{I}$ only depends on the graphical representation in the time interval $[t-1,t]$.
@@
$\begin{split}\mathbb{P}\big{(}\mathcal{B}_{\tau}\mid\mathcal{F}_{\tau}\big{)}&\geq\mathbb{P}\left(\inf_{2\leq k\leq\bar{k}}|\mathcal{G}_{\rho,k}^{\prime}|\geq\delta\lambda TNp(N,L)\right)\\ &\geq\mathbb{P}\left(\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}\Big{(}\inf_{2\leq k\leq\bar{k}}|\mathcal{G}_{\rho,k}^{\prime}|\leq\delta\lambda TNp(N,L)\,\Big{|}\,\mathcal{F}_{0,S}(x_{1})\Big{)}\mathds{1}_{\mathcal{S}_{\rho,N}(x_{1})}\right)\\ &\geq(1-R_{\lambda}(N))\mathbb{P}(\mathcal{S}_{\rho,N}(x_{1}))\geq(1-R_{\lambda}(N))b_{2}.\end{split}$
@@
where $c:=2^{-1}b_{1}^{2}b_{2}^{2}b_{3}\mathbb{P}(\zeta<L)$. Under the assumption that $r=r_{N}$ satisfies Condition ($\star$7.11), we have established the following lower bound for $g$,
@@
The following lemma tells us that under the Condition (3.7) we can find a natural number $r_{N}$ such that Condition ($\star$7.11) holds for $N$ large enough.
@@
where $c$ is chosen as in Lemma 7.11. Then there exists a subsequence $(N_{n})_{n\geq 0}$ with $N_{n}\to\infty$ as $n\to\infty$ such that $r_{N_{n}}\to\infty$ as $n\to\infty$ and Condition (italic-$\star$7.11) is satisfied for $n$ large enough.
@@
For the rest of this proof we write $T_{n}=T_{N_{n}}$ and $r_{n}=r_{N_{n}}$. Note that for this sequence it also holds that $N_{n}\mathbb{P}(\zeta=N_{n})\to 0$ as $n\to\infty$, which implies that $r_{n}\to\infty$ as $n\to\infty$.
@@
where we can avoid the ceiling and the floor in the factors $S(8r_{n}+4T_{n})^{-1}$ and $\bar{k}$ as $\lceil S(8r_{n}+4T_{n})\rceil/S(8r_{n}+4T_{n})\to 1$ as $n\to\infty$ (and similarly for $\bar{k}$ and $r_{n}$). Therefore, (7.32) is equivalent to ($\star$7.11) for $N_{n}$ large enough. Now, we see that (7.32) is equivalent to
@@
Suppose $N,\lambda_{N},\eta,\alpha$ satisfy (R’) and assume that Condition (3.8) holds. Then $r=r_{N}$ defined as in (7.30) satisfies Condition (italic-$\star$7.11) for some large $N$.
@@
guarantees together with (7.39) that (7.41) holds for $N_{n}$ large enough since the limit diverges. In other words, $r=r_{N_{n}}$ defined as in (7.30) satisfies Condition ($\star$7.11) for some large $N_{n}$. ∎
@@
Let us begin with the first claim. Assume that Condition (3.7) holds, $\alpha<1$ and $\eta\in(-\infty,\tfrac{1-\alpha}{2})$. Fix $\lambda\in\big{(}0,\tfrac{1}{2}\big{)}$. Next define $r_{N}$ as in (7.30). By Lemma 7.12, we know $r_{N}$ satisfies Condition ($\star$7.11) for some $N$ large enough. Now, we are able to apply Lemma 7.11 to deduce that the non-negative and non-decreasing function

### v2

@@ 7.4 Transmitting the infection from one star to another
$\mathbb{P}_{\mathcal{T},N}(B_{m}^{1}\mid\mathcal{F}_{t_{m},S})\mathds{1}_{D_{m}}\geq\mathbb{P}_{\mathcal{T},N}(B_{m}^{1}\mid\mathcal{F}_{t_{m},S}\,,\,\rho\notin\mathbf{C}^{*-x}_{t_{m}})\mathds{1}_{D_{m}}\geq\frac{\lambda\lfloor\delta\lambda TNp(N,L)\rfloor T}{(\lambda\lfloor\delta\lambda TNp(N,L)\rfloor+1)T+1}\mathds{1}_{D_{m}}.$
@@
where in the last step we treated the first time interval $0,t_{2}]$ differently, since we know that at time 0 the root $\rho$ is infected, and thus we do not need to reinfect it in the first time period. Recalling that we chose $m=\lfloor S(8r+4T)^{-1}\rfloor$, we can conclude
@@
$\leq\mathbb{P}_{\mathcal{T},N}\bigg(\bigcap_{i=1}^{m}A_{i}\cap\bigcap_{j=2}^{m}D_{j}\Big|\mathcal{F}_{0,S}\bigg)\leq\bigg(1-b_{\lambda}(N)C_{p}(N)\bigg(\frac{\lambda\nu_{1}\kappa_{1}c_{p}}{\lambda+\lambda\nu_{1}+\nu_{1}+1}\bigg)^{r}\bigg)^{\lfloor S(8r+4T)^{-1}\rfloor},$
@@
We recall that the first claim of Theorem 3.4 is that if $\alpha\in[0,1)$, $\eta\in(-\infty,\tfrac{1-\alpha}{2})$ and (3.7) holds then $\lambda_{2}=0$. In other words, for every $\lambda>0$ strong survival takes place with positive probability. The second statement is that if $\alpha\in[0,1)$ and (3.8) holds then $\lambda_{2}<\infty$, i.e. for $\lambda$ large enough strong survival takes place with positive probability. Note that it is sufficient to consider $\eta\geq 0$ in order to show this claim because the case $\eta<0$ is covered by the first claim. Therefore, we need to consider two parameter regimes. First let
@@
where $C$ is the universal constant given in Lemma 7.2. Note that if we show that there is strong survival for all $\lambda\in\big(0,\tfrac{1}{2}\big)$, then it follows by monotonicity for all $\lambda>0$. The second parameter regime is the following
$\alpha\in[0,1),\,\eta\geq 0,\,N\in\mathbb{N},\,\lambda=\lambda_{N}=(3T_{N})^{-1}\text{ such that }N>C\vee L.$ (R’)
Observe that in both cases we have that $2\lambda T_{N}<1$ for all $N$ large enough, which is necessary for several results in the previous sections.
We recall from Section 3.2 that our strategy is to apply [18, Lemma 6.2] to the function $g(t):=\mathbb{P}^{\{\rho\}}_{N}(\rho\in\mathbf{C}_{t}),\ t\geq 0.$ In order to do so, we need to find a non-decreasing function $H$ on $0,\infty)$ such that $H(x)\geq x$ in a neighbourhood of 0 and show that $g$ and $H$ satisfy
@@
where $S$ is chosen as in (7.17). The next lemma will be useful in our proof, its proof can be found in [18, Lemma 2.3].
Lemma 7.11.
Let $M$ be a positive integer-valued random variable and pick $p<\mathbb{E}[M]$. For any $x>0$, let $M_{x}$ be a $\text{Bin}(M,x)$ random variable. Then there exists an $\epsilon>0$ such that $\mathbb{P}(M_{x}\geq 1)\geq px\wedge\epsilon$.
For simplicity of exposition, we split the proof of Theorem 3.4 into several lemmas. The first lemma gives us a good candidate for the function $H$.
@@
There exists $c\in(0,1]$ and $\epsilon>0$ such that for $N$ large enough, the function
@@
$p_{in}:=\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}\big(x\in\mathbf{C}_{s}\ \text{for some}\ s\in[0,S]\mid\mathcal{F}_{0,S}\big).$
Recall that the process $(\mathbf{C}_{t}^{*-x})_{t\geq 0}$ is defined via the same graphical representation as $(\mathbf{C}_{t})_{t\geq 0}$ but the infection is restricted to $\mathcal{V}^{*-x}=\mathcal{N}_{\rho}\cup\{\rho,x_{1},\dots,x_{r-1},x\}$. We begin by noting that due to the monotonicity of the CPDG, we have
$p_{in}\geq\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}\big(x\in\mathbf{C}_{s}^{*-x}\ \text{for some}\ s\in[0,S]\mid\mathcal{F}_{0,S}\big).$
Now, Lemma 7.10 and monotonicity tells us that
$\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}\big(x\in\mathbf{C}_{s}^{*-x}\ \text{for some}\ s\in[0,S]\mid\mathcal{F}_{0,S}\big)\geq 1-(F_{\lambda}(N,r)+R_{\lambda}(N)),$
on the event $\mathcal{S}_{\rho,N}$, where $S$ is defined as in Lemma 7.8, the functions $R_{\lambda}(N)$ and $F_{\lambda}(N,r)$ are defined in (7.18) and (7.19), respectively. Then, using the inequality $1-x\leq e^{-x}$, observe that ($\star$7.12) forces $F_{\lambda}(N,r)$ to be at most $e^{-4}$, i.e.
@@
Before we proceed we need to introduce some further notation. Given a realisation of the underlying tree $\mathcal{T}$ we denote by $\mathcal{S}_{x,N}^{t-S}$ the event that $x\in\mathbf{P}_{r}$ is a stable star starting from time tSt-S, where $t\geq S$. Furthermore, we explicitly exclude the neighbour in generation r1r-1, but we also omit this in the notation to avoid clutter. In the following, we will condition on the event
@@
i.e. that $x\in\mathbf{C}_{t-S}$ for some $x\in\mathbf{P}_{r}$ which is a stable star starting at time tSt-S. We deduce for $t>S$
@@
Then, by using the fact that $\mathbf{C}_{s}^{*-x}\subset\mathbf{C}_{s}$ for all $s\geq 0$, followed by the tower property, (7.24) and Remark 7.9, we deduce
$\mathbb{E}_{N}\big[M_{r}^{S}\big]\geq\mathbb{E}_{N}\left[\sum_{x\in\mathbf{P}_{r}}\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}(x\in\mathbf{C}_{s}^{*-x}\ \text{for some}\ s\in[0,S]\mid\mathcal{F}_{0,S})\mathds{1}_{\mathcal{S}_{\rho,N}}\right]\geq b_{1}\mathbb{E}_{N}\bigg[\sum_{x\in\mathbf{P}_{r}}\mathds{1}_{\mathcal{S}_{\rho,N}}\bigg].$
@@
$\mathbb{E}_{N}\bigg[\sum_{x\in\mathbf{P}_{r}}\mathds{1}_{\mathcal{S}_{\rho,N}}\bigg]$ $\geq\mathbb{E}_{N}\big[|\mathbf{P}_{r}|\big]\mathbb{P}_{N}(\mathcal{S}_{\rho,N})\geq\mathbb{P}(\zeta<L)N\mu^{r-1}_{L}\mathbb{P}(\zeta=N)\mathbb{P}_{N}(\mathcal{S}_{\rho,N}),$
In addition, by Lemma 7.1 we have that $\mathbb{P}_{N}(\mathcal{S}_{\rho,N})\geq 1-e^{-c_{L}Np(N,L)}$. Thus there exists $b_{2}>0$ such that for $N$ large enough $\mathbb{P}_{N}(\mathcal{S}_{\rho,N})\geq b_{2}$. In other words, we deduce
@@
Now, ignore all the infections of $x\in\mathbf{P}_{r}$ by its parent except the first infection. Then, the contact process on the subtrees rooted at vertices $x\in\mathbf{P}_{r}$ which are infected at some time $s<S$ will evolve independently from time $s$ to time tSt-S, since by the branching property of the BGW tree and using that the background process is stationary. Then the vertex $x$ will be infected at time tSt-S with probability at least $\chi(t)$.
Furthermore, the event that $x\in\mathbf{P}_{r}$ is a stable star from tSt-S onwards, i.e. that $\mathcal{S}_{x,N}^{t-S}$ is satisfied, is independent of everything that happened before time tSt-S and by symmetry it follows that $b_{2}$ is a lower bound for the probability of $x\in\mathbf{P}_{r}$ being a stable star for every realisation $\mathcal{T}$, i.e. $\mathbb{P}_{\mathcal{T},N}(\mathcal{S}^{t-S}_{x,N})\geq b_{2}$.
If we denote by $\widetilde{M}_{r}^{t-S}$ the random number of vertices in $\mathbf{P}_{r}$ which are infected at time tSt-S and are stable stars starting at tSt-S, we can conclude that $\widetilde{M}_{r}^{t-S}$ stochastically dominates a random variable $M_{\chi}$ that has distribution $\text{Bin}(M_{r}^{S},b_{2}\chi(t))$. Appealing to Lemma 7.11 there exists $\epsilon_{1}>0$ such that
@@
For that purpose we need to introduce a bit more notation. Let $\mathcal{F}_{s}$ be the $\sigma$-algebra which is generated by the graphical construction up to time $s$. Furthermore, we denote by $\mathcal{F}^{r}$ the $\sigma$-algebra generated by $\Delta^{\text{rec}}_{z}$ and $\Delta^{\text{up}}_{\{y,z\}}$ for all $z\in\mathcal{V}_{r+1}$, i.e. vertices in generation $r$, and $\{y,z\}\in\mathcal{E}$ with $d(\rho,y)=r$, i.e. edges between generation $r$ and $r+1$. In words this means that $\mathcal{F}^{r}$ contains all the information about recoveries in generation $r+1$ and update events between generation $r$ and $r+1$ for the whole time period 0 to $\infty$. Finally we define the $\sigma$-algebra $\mathcal{F}_{s}^{r}:=\sigma(\mathcal{F}_{s}\cup\mathcal{F}^{r})$ for $s\geq 0$.
@@
Recall that on the event $A_{t-S}$ we know that there exists an $x\in\mathbf{P}_{r}$ which is a stable star from tSt-S onwards and $x\in\mathbf{C}_{t-S}$, i.e. $x$ is infected at time tSt-S. This means we are again in a situation that we can consider the restriction of the graphical representation to a path of length $r$ between $\rho$ and the neighbourhood of $x$. Then, it follows analogously as in the proof of Lemma 7.10 and (7.24) that
@@
where we used that the path is between generation 0 and $r$, and thus independent of everything in generation $r+1$ and higher. Therefore, we can conclude that $h(t)\geq b_{1}$.
@@
Note that on $\mathcal{B}_{\tau}$ we have at least $\delta\lambda TNp(N,L)$ good neighbours in $\tau+4T,t-1]$. Further, define $\mathcal{I}$ to be the event that at least one of the $\delta\lambda TNp(N,L)$ neighbours which are infected at time t1t-1 infects $\rho$ at a time in [t1,t][t-1,t] before this neighbour recovers and
@@
Now we find a lower bound for the two conditional probabilities on the right-hand side of the previous equality. First note that $\mathbb{P}(\mathcal{R}_{\rho}\ |\mathcal{F}_{t-1})\geq e^{-1}$. Recall that $\mathcal{N}_{\rho}$ are the neighbours of $\rho$ with degree less than $L$, i.e. $D_{x_{i}}\leq L$ for all $i\in\{1,\dots,\tilde{N}\}$. Denote $\tilde{N}:=\lfloor\delta\lambda TNp(N,L)\rfloor$ and by $x_{1},\dots,x_{\tilde{N}}\in\mathcal{N}_{\rho}$ the $\tilde{N}$ neighbours that are infected at time t1t-1. Note that we can rewrite the event $\mathcal{I}$ as follows
@@
where we used that $\mathcal{I}$ only depends on the graphical representation in the time interval [t1,t][t-1,t].
@@
$\begin{split}\mathbb{P}\big(\mathcal{B}_{\tau}\mid\mathcal{F}_{\tau}\big)&\geq\mathbb{P}\left(\inf_{2\leq k\leq\bar{k}}|\mathcal{G}_{\rho,k}^{\prime}|\geq\delta\lambda TNp(N,L)\right)\\ &\geq\mathbb{P}\left(\mathbb{P}_{\mathcal{T},N}^{\{\rho\}}\Big(\inf_{2\leq k\leq\bar{k}}|\mathcal{G}_{\rho,k}^{\prime}|\leq\delta\lambda TNp(N,L)\,\Big|\,\mathcal{F}_{0,S}\Big)\mathds{1}_{\mathcal{S}_{\rho,N}}\right)\\ &\geq(1-R_{\lambda}(N))\mathbb{P}(\mathcal{S}_{\rho,N})\geq(1-R_{\lambda}(N))b_{2}.\end{split}$
@@
where $c:=2^{-1}b_{1}^{2}b_{2}^{2}b_{3}\mathbb{P}(\zeta<L)$. Under the assumption that $r=r_{N}$ satisfies Condition ($\star$7.12), we have established the following lower bound for $g$,
@@
The following lemma tells us that under the Condition (3.7) we can find a natural number $r_{N}$ such that Condition ($\star$7.12) holds for $N$ large enough.
@@
where $c$ is chosen as in Lemma 7.12. Then there exists a subsequence $(N_{n})_{n\geq 0}$ with $N_{n}\to\infty$ as $n\to\infty$ such that $r_{N_{n}}\to\infty$ as $n\to\infty$ and Condition ($\star$7.12) is satisfied for $n$ large enough.
@@
For the rest of this proof we write $T_{n}=T_{N_{n}}$ and $r_{n}=r_{N_{n}}$. Note that for this sequence it holds that $N_{n}\mathbb{P}(\zeta=N_{n})\to 0$ as $n\to\infty$, which implies that $r_{n}\to\infty$ as $n\to\infty$.
@@
where we can avoid the ceiling and the floor in the factors $S(8r_{n}+4T_{n})^{-1}$ and $\bar{k}$ as $\lceil S(8r_{n}+4T_{n})\rceil/S(8r_{n}+4T_{n})\to 1$ as $n\to\infty$ (and similarly for $\bar{k}$ and $r_{n}$). Therefore, (7.31) is equivalent to ($\star$7.12) for $N_{n}$ large enough. Now, we see that (7.31) is equivalent to
@@
Suppose $N,\lambda_{N},\eta,\alpha$ satisfy (R’) and assume that Condition (3.8) holds. Then $r=r_{N}$ defined as in (7.29) satisfies Condition ($\star$7.12) for some large $N$.
@@
guarantees together with (7.38) that (7.40) holds for $N_{n}$ large enough since the limit diverges. In other words, $r=r_{N_{n}}$ defined as in (7.29) satisfies Condition ($\star$7.12) for some large $N_{n}$. ∎
@@
Let us begin with the first claim. Assume that Condition (3.7) holds, $\alpha<1$ and $\eta\in(-\infty,\tfrac{1-\alpha}{2})$. Fix $\lambda\in\big(0,\tfrac{1}{2}\big)$. Next define $r_{N}$ as in (7.29). By Lemma 7.13, we know $r_{N}$ satisfies Condition ($\star$7.12) for some $N$ large enough. Now, we are able to apply Lemma 7.12 to deduce that the non-negative and non-decreasing function